OpenStudy (anonymous):

Solve the differential equation. Find the particular solution that goes through the point (0,4). dy/dx = (x)/(x^2 -9)

OpenStudy (anonymous):

\[\int\limits_{}^{}dy=\int\limits_{}^{}\frac{ x }{ x^2-9 }*dx\]

OpenStudy (anonymous):

let x^2-9=t 2x*dx=dt x*dx=dt/2

OpenStudy (anonymous):

\[y=\int\limits_{}^{}\frac{ dt }{ 2t }\]

OpenStudy (anonymous):

so do u sub???

OpenStudy (gorv):

yeah substitution

OpenStudy (gorv):

+ c will also be their

OpenStudy (gorv):

\[y=\int\limits_{}^{}\frac{ dt }{ 2*t } +c\]

OpenStudy (anonymous):

take the integral and then substitute for u??

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

yeah right

OpenStudy (anonymous):

@gorv i forgot to add c loll

OpenStudy (anonymous):

so du^2 /4u +c?

OpenStudy (anonymous):

@g17

OpenStudy (anonymous):

how u got it loll??? tell me

OpenStudy (anonymous):

wait what? when you take the integral of du/2u its du^2/4u +c or no??/

OpenStudy (anonymous):

@gorv when you take the integral of du/2u its du^2/4u +c or no??/

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ du }{ 2u }+c=\frac{ 1 }{ 2 }*\int\limits_{}^{}\frac{ du }{ u }+c=\frac{ 1 }{ 2 }*Ln(u)+c\]

OpenStudy (anonymous):

would the Ln be in absolute value since u could be positive or negative @g17

OpenStudy (anonymous):

how u got du^2...can u tell me

OpenStudy (anonymous):

ln is natural log

OpenStudy (anonymous):

so take positive value

OpenStudy (anonymous):

@pnp97

OpenStudy (anonymous):

i know ln is natural log but when it can be negative dont u put it in absolute value and umm idk i was taking an integral but i forgot that you could take out the 1/2

OpenStudy (anonymous):

oh ok so it has to be positive

OpenStudy (anonymous):

so now you plug in x

OpenStudy (anonymous):

x^2 -9 for u @g17

OpenStudy (anonymous):

yeah right

OpenStudy (anonymous):

and then is it 4= 1/2 ln(0^2 -9) +c?

OpenStudy (anonymous):

@g17

OpenStudy (anonymous):

yeah take absolute or positive value

OpenStudy (anonymous):

as negative value is not allowed in log

OpenStudy (anonymous):

so ln absv(0^2- 9)?

OpenStudy (anonymous):

@g17

OpenStudy (anonymous):

yeah u r right

OpenStudy (anonymous):

do you solve for c now? @g17

OpenStudy (anonymous):

yeah and out value of c in equation that will be your particular solution

OpenStudy (anonymous):

@pnp97

OpenStudy (anonymous):

ohh ok! so 4= 1/2ln absv(-9) +c @g17 how do you further simplify it?

OpenStudy (anonymous):

c=4-1/2 ln9 y=1/2lnu+4-1/2ln9 y=1/2(lnu-ln9)+4 y=1/2ln(u/9)+4 that's it

OpenStudy (anonymous):

thanks so much!:-) @g17